If a 5 kg object experiences a 10 N force for a duration of 0.1 s, then what is the change in momentum of the object?


1
Expert's answer
2022-01-10T09:09:41-0500

According to the second Newton's law, the change in momentum is equal to:


"\Delta p = F\Delta t"

where "F = 10N" is the net force, and "\Delta t = 0.1s" is the duration. Thus, obtain:


"\Delta p = 10N\cdot 0.1s = 1kg\cdot m/s"

Answer. "1kg\cdot m/s".